# Neutral theory of cooperators

### v1

@@ 2 Steady-state species abundance distribution
$=\left(\frac{1-\mu}{N^{3}}\right)^{n-1}(n-1)!\frac{\left(N-1\right)!}{\left(N-n\right)!}\frac{\left(N-1\right)!}{\left(N-n\right)!}$ (S2.12)
@@
$=\frac{\big{(}\frac{1-\mu}{1-\lambda^{*}_{\circ}}\big{)}^{n}}{n(N-n)}\frac{(1-\lambda^{*}_{\circ})^{n}{\lambda^{*}_{\circ}}^{-n}}{\left(N-n\right)!\,\Gamma\left(N/\lambda^{*}_{\circ}-N+n+1\right)}$ (S2.15)
@@
$\lambda^{*}_{\circ}\approx{\lambda^{*}}:=\langle\lambda(\mathbf{n})\rangle$ (S2.20)
[…]
To justify (S2.20), let us consider the definitions (S1.2)-(S1.3), and observe that
$\lambda(\mathbf{n})$ $=\left(\frac{N}{N-n}\right)^{2}\lambda(\mathbf{n}^{\circ})+\left(\frac{n}{N}\right)^{2}\geq\lambda(\mathbf{n}^{\circ}).$ (S2.21)
For the upper bound, we may rewrite the equation as
$\lambda(\mathbf{n})$ $=\lambda(\mathbf{n}^{\circ})+\frac{2Nn-n^{2}}{(N-n)^{2}}\lambda(\mathbf{n}^{\circ})+\left(\frac{n}{N}\right)^{2}$
$\leq\lambda(\mathbf{n}^{\circ})+\frac{2n/N}{(1-n/N)^{2}}+\left(\frac{n}{N}\right)^{2}$ (S2.22)
where in the last line we used that $n^{2}\geq 0$ and $\lambda(\mathbf{n}^{\circ})\leq 1$. Observe that $\langle n\rangle/N$ is the inverse of the number of non-extinct species, see Eq. (20) in the main text. Thus, assuming that there is a large number of species in the system, the steady-state probability distribution should be concentrated on $n\ll N$. Then, taking expectations of both sides and ignoring terms of order $n/N$ gives ${\lambda^{*}}:=\langle\lambda(\mathbf{n})\rangle\approx\lambda^{*}_{\circ}:=\langle\lambda_{\circ}(\mathbf{n})\rangle$.
@@ 3 Steady-state Simpson index
$\lambda(\mathbf{n})\approx\sum_{n=1}^{N}r_{n}\left(\frac{n}{N}\right)^{2}\,$ (S3.23)
@@
$\frac{\langle r_{n}\rangle n}{N}=\frac{P_{n}n}{\sum_{n^{\prime}}P_{n^{\prime}}n^{\prime}}.$ (S3.24)
@@
${\lambda^{*}}=\langle\lambda(\mathbf{n})\rangle\approx\sum_{n=1}^{N}\left(\frac{\langle r_{n}\rangle n}{N}\right)\left(\frac{n}{N}\right)=\frac{1}{N}\frac{\sum_{n=1}^{N}P_{n}n^{2}}{\sum_{n^{\prime}=1}^{N}P_{n^{\prime}}n^{\prime}},$ (S3.25)
@@ 3.1 Low migration Simpson index
[…]
@@
[…]
@@ 3.2 High migration Simpson index derived from (12)
as follows by considering the first two terms of the continued fraction derived by Laplace [49, Livre X, p. 255]. Plugging into (12) and solving gives
${\lambda^{*}}\approx\frac{\mu-\sqrt{\mu^{2}-4/N}}{2}\approx\frac{1}{N\mu}$ (S3.31)
@@ 4 Steady-state number of species
$\frac{N_{\mathsf{out}}}{N}\approx\Phi\left(\frac{1/{\lambda^{*}}-N{\lambda^{*}}}{\sqrt{N}}\right)$ (S4.34)
@@
$\sum_{n=1}^{1/{\lambda^{*}}}\frac{1}{n}e^{-(n-N{\lambda^{*}})^{2}/2N}\approx e^{-(1-N{\lambda^{*}})^{2}/2N}\sum_{n=1}^{1/{\lambda^{*}}}\frac{1}{n}\approx-\sqrt{2\pi N}\mu\ln{\lambda^{*}},$ (S4.38)
where in the last approximation we evaluated the (slow-changing) exponential term in the range $1\leq n\leq 1/{\lambda^{*}}$ as $e^{-(1-N{\lambda^{*}})^{2}/2N}\approx e^{N{\lambda^{*}}^{2}/2}$, and then plugged in our estimation of ${\lambda^{*}}$ in the low-migration regime (Eq. (12) in the main text). Combining the above gives (25) as reported in the main text.
@@ 5.1 Maximum abundances distribution
As described above, our model can be reduced to a birth-death process with transition probabilities $b_{n}$ and $d_{n}$, Eqs. (S1.7)-(S1.8). We are also using the approximation ${\lambda^{*}}(\mathbf{n})\approx{\lambda^{*}}$. Then, the probability $u_{n}(\ell)$ must obey two boundary conditions:
@@ 5.2 Bimodality of $Q_{m}$
$0=\frac{d}{dm}Q_{m}\approx-\frac{d^{2}}{dm^{2}}u_{1}(m)\propto-2\left(\frac{d}{dm}v_{m}\right)^{2}+v(m)\frac{d^{2}}{dm^{2}}v_{m}$ (S5.50)
where, for notational convenience, we introduced $v_{m}=1/u_{1}(m)$ to denote the right side of Eq. (S5.48). Considering that equation, it can be shown that the derivatives of $v_{m}$ become very small at $m=N{\lambda^{*}}$, where the minimum of the exponent’s argument is reached. Linearizing the right side of Eq. (S5.50) around $m\approx N{\lambda^{*}}$ and then solving gives
@@
In Eq. (S5.56), we show that $\sqrt{2N}\mathscr{D}(x_{1})\approx 1/{\lambda^{*}}$. Since ${\lambda^{*}}\gg 1/\sqrt{N}$ in the low-migration regime, the denominator in the last term in Eq. (S5.51) is exponentially large, allowing us to ignore this whole term. We also have $N{\lambda^{*}}\gg 1/2$, thus dropping all sub-dominant terms gives the approximation
@@ 5.3 Infiltration probability
Here we derive the infiltration probability to the dynamical core, $\beta$. We will use the asymptotic expansion of the Dawson function [50, 42:6:4],
@@ 6 Residence times
In this section, we calculate the mean residence time of species whose trajectories reach the dynamical core, as defined in SM5. We also argue that the cooperator core obtained in our study of the steady-state abundance distribution and the dynamical core are effectively equivalent.
At stationarity, let us denote by $T_{n}$ the mean first passage time (MFPT) from a species with initial abundance $n$ to reach extinction (at abundance 00). We observe that a core species will spend a long time at high abundance values; hence, we assume that the mean residence of a species whose trajectory belongs to the dynamical core can be approximated by the MFPT from $N{\lambda^{*}}$ to extinction,
@@
$v_{n}=U_{n-1}v_{1}-Q_{n-1},\quad\text{with}\quad U_{m-1}:=\prod_{k=1}^{m-1}\frac{d_{k}}{b_{k}}\quad\text{and}\quad Q_{m-1}=\sum_{k=1}^{m-1}\frac{1}{b_{k}}\prod_{j=k+1}^{m-1}\frac{d_{j}}{b_{j}}.$ (S6.60)
@@
$Q_{m-1}=\sum_{k=1}^{m-1}\frac{1}{b_{k}}\prod_{j=k+1}^{m-1}\frac{d_{j}}{b_{j}}=U_{m-1}\sum_{k=1}^{m-1}\frac{1}{b_{k}}\prod_{j=1}^{k}\frac{b_{j}}{d_{j}}.$ (S6.62)
The expression for $Q_{m-1}$ is related to the steady-state distribution $P_{k}$ because
@@
therefore we may rewrite $Q_{m-1}$ as
$Q_{m-1}=\frac{U_{m-1}}{P_{1}d_{1}}\sum_{k=1}^{m-1}P_{k}=\frac{U_{m-1}}{P_{1}d_{1}}C(m-1).$ (S6.64)
@@
The expected number of species in steady-state, $R^{*}$, satisfies the balance equation between outflow due to extinction and inflow due to migration:
@@
Next, we observe that in the low-migration regime, $P_{k}$ is well-approximated as a Gaussian distribution with mean $N{\lambda^{*}}$ and standard deviation $\sqrt{N}$ (see Eq. (10) in the main text). Hence, $C$ is approximated as the cumulative density function of this Gaussian distribution, so $1-C(m-1)\approx 1$ for small $m$ and $21-C(m-1)\approx 1/$ 2 for $m\approx N{\lambda^{*}}$. On the other hand, for our range of $m\in[1,N{\lambda^{*}}]\cap\mathbb{N}$, $U_{m-1}$ has a local maximum at $m=1$ and, at small values of $m$, decays exponentially with rate $1/{\lambda^{*}}$, see Eq. (S6.61). Therefore, in the low-migration regime where $N{\lambda^{*}}^{2}\gg 1$, one can verify that as $m$ increases, $U_{m-1}$ decays essentially to 0 while $1-C(m-1)$ remains very close to 1. This allows us to approximate:
@@
The right-hand side corresponds to the expression for $1/u_{1}(N{\lambda^{*}})$ that we approximated in (S5.48). Using our definition of $\beta$ (S5.54) and our result (S5.55) in terms of the Dawson function, we combine with (S6.71) to give
$T_{N{\lambda^{*}}}\approx\frac{R^{*}}{\mu}\sqrt{2N}\mathscr{F}_{N}\left(\frac{N{\lambda^{*}}-1}{\sqrt{2N}}\right)\approx\frac{R^{*}}{\mu\beta}.$ (S6.73)

### v2

@@ 2 Steady-state species abundance distribution
$=\left(\frac{1-\mu}{N^{3}}\right)^{n-1}(n-1)!\frac{\left(N-1\right)!}{\left(N-n\right)!}\frac{\left(N-1\right)!}{\left(N-n\right)!}\,.$ (S2.11)
@@
$=\frac{\big(\frac{1-\mu}{1-\lambda^{*}_{\circ}}\big)^{n}}{n(N-n)}\frac{(1-\lambda^{*}_{\circ})^{n}{\lambda^{*}_{\circ}}^{-n}}{\left(N-n\right)!\,\Gamma\left(N/\lambda^{*}_{\circ}-N+n+1\right)}\,.$ (S2.14)
@@
$\lambda^{*}_{\circ}\approx{\lambda^{*}}:=\langle\lambda(\mathbf{n})\rangle\,.$ (S2.19)
where $\langle\cdot\rangle$ indicates expectations under the steady-state distribution of abundances $\mathbf{n}$. Combining the above approximations, we are left with the probability distribution reported in Eq. (10) in the main text.
2.1 Quality of the approximation (S2.19)
Eq. (S2.19), a key step in the derivation above, involves two approximations. First, we approximate the Simpson index of the non-representative species $\lambda_{\circ}(\mathbf{n})$ with the Simpson index of all species $\lambda(\mathbf{n})$. Second, we ignore steady-state fluctuations of the Simpson index, approximating $\lambda(\mathbf{n})\approx{\lambda^{*}}:=\langle\lambda\rangle$. In this section, we study the error involved in both approximations.
To consider the approximation $\lambda_{\circ}(\mathbf{n})\approx\lambda(\mathbf{n})$, we use the identity
$N^{2}\lambda={(N-n)^{2}\lambda_{\circ}+n^{2}}\,$, (S2.20)
where we drop the arguments in $\lambda(\mathbf{n})$ and $\lambda_{\circ}(\mathbf{n})$ for simplicity of notation. It can be rearranged as
$N^{2}(\lambda_{\circ}-\lambda)$ $=2Nn\lambda_{\circ}-n^{2}(\lambda_{\circ}+1)$
$(N^{2}-2Nn)(\lambda_{\circ}-\lambda)$ $=2Nn\lambda-n^{2}(\lambda_{\circ}+1)\,.$
Dividing both sides by $N^{2}\lambda$ gives
$\left(1-\frac{2n}{N}\right)\frac{\lambda_{\circ}-\lambda}{\lambda}=\frac{2n}{N}-\frac{n^{2}}{N^{2}}\frac{\lambda_{\circ}+1}{\lambda}\,.$
Next, we drop small terms $2n/N\ll 1$ and $\lambda_{\circ}\ll 1$ to give
$\frac{\lambda_{\circ}-\lambda}{\lambda}\approx\frac{2n}{N}-\frac{n^{2}}{N^{2}}\frac{1}{\lambda}\,.$
The results described in the main text (and rest of this SI Appendix) imply that steady-state abundances are concentrated around $n\approx N\lambda^{*}$. Plugging this in and simplifying gives
$\frac{\lambda_{\circ}-\lambda}{\lambda}\approx{\lambda^{*}}\approx\sqrt{\frac{-\ln(2\pi N\mu^{2})}{N}}\,$, (S2.21)
where we used Eq. (13). For instance, for the system in Figure 3(b) ($N=10^{5},\mu=10^{-5}$), the relative error is $\approx 0.01$.
Next, we consider the approximation $\lambda(\mathbf{n})\approx{\lambda^{*}}:=\langle\lambda\rangle$ that ignores fluctuations. To estimate the scale of the fluctuations of the Simpson index, we use the definition
$\lambda(\mathbf{n}):=\sum_{i=1}^{R}({n_{i}}/{N})^{2}\,$, (S2.22)
where $R$ is the number of species and $n_{i}$ is abundance of species $i$. In the steady state of the low-migration regime, each $n_{i}$ may be approximated by a Gaussian with mean $N{\lambda^{*}}$ and variance $N$. Hence $n_{i}/N$ may be approximated as a Gaussian with mean $\lambda^{*}$} and variance $1/N$, while $(n_{i}/N)^{2}$ may be approximated as the square of this Gaussian, which has variance $4{\lambda^{*}}^{2}/N+2/N^{2}\approx 4{\lambda^{*}}^{2}/N$. Assuming that steady-state abundance fluctuations are approximately independent across species, the fluctuations of $\lambda(\mathbf{n})$ will have variance $4R^{*}{\lambda^{*}}^{2}/N$.
The scale of relative fluctuations of $\lambda(\mathbf{n})$ are given by the ratio of the standard deviation and the mean, ${\textrm{Std}(\lambda)}/{{\lambda^{*}}}\approx\sqrt{4R^{*}/N}$. But recall from Eq. (20) that $R^{*}=N/\langle n\rangle\approx 1/{\lambda^{*}}$. Combining and using Eq. (S2.21) gives a simple estimate of the relative fluctuations:
$\frac{\textrm{Std}(\lambda)}{{\lambda^{*}}}\approx\frac{2}{\sqrt{{\lambda^{*}}N}}\approx 2^{3/4}(-\ln\mu)^{-1/4}N^{-1/4}\,.$ (S2.23)
For instance, for the system shown in Figure 3(b), $N=10^{5}$ and $\mu=10^{-5}$ and the size of relative fluctuations is $\approx 0.05$.
@@ 3 Steady-state Simpson index
$\lambda(\mathbf{n})=\sum_{n=1}^{N}r_{n}\left(\frac{n}{N}\right)^{2}\,$ (S3.24)
@@
$\frac{\langle r_{n}\rangle n}{N}=\frac{P_{n}n}{\sum_{n}P_{n}n}.$ (S3.25)
@@
${\lambda^{*}}=\langle\lambda(\mathbf{n})\rangle\approx\sum_{n=1}^{N}\left(\frac{\langle r_{n}\rangle n}{N}\right)\left(\frac{n}{N}\right)=\frac{1}{N}\frac{\sum_{n=1}^{N}P_{n}n^{2}}{\sum_{n=1}^{N}P_{n}n},$ (S3.26)
@@ 3.1 Low migration Simpson index
${\lambda^{*}}\approx\frac{1}{N}\frac{\int_{1}^{N}e^{-\left[n-N({{\lambda^{*}}}-\mu)\right]^{2}/2N}n\,dn}{\int_{1}^{N}e^{-\left[n-N({{\lambda^{*}}}-\mu)\right]^{2}/2N}dn}\,.$ (S3.28)
@@
$\frac{1}{N}\int_{1}^{N}e^{-\left[n-N({{\lambda^{*}}}-\mu)\right]^{2}/2N}n\,dn$ $\approx e^{-N({\lambda^{*}}-\mu)^{2}/2}+\sqrt{\frac{\pi N}{2}}({\lambda^{*}}-\mu)\ \textrm{erfc}\left(\sqrt{\frac{N}{2}}(\mu-{\lambda^{*}})\right)$
$\int_{1}^{N}e^{-\left[n-N({{\lambda^{*}}}-\mu)\right]^{2}/2N}dn$ $\approx\sqrt{\frac{\pi N}{2}}\textrm{erfc}\left(\sqrt{\frac{N}{2}}(\mu-{\lambda^{*}})\right),$ (S3.29)
where we used the complementary error function, $\textrm{erfc}(z):=2/\sqrt{\pi}\int_{z}^{\infty}e^{-t^{2}}dt$. Combining with (S3.28) yields
@@ 3.2 High migration Simpson index derived from (12)
as follows by considering the first two terms of the continued fraction derived by Laplace [58, Livre X, p. 255]. Plugging into (12) and solving gives
${\lambda^{*}}\approx\frac{\mu-\sqrt{\mu^{2}-4/N}}{2}\approx\frac{1}{N\mu}\,$, (S3.32)
@@ 4 Steady-state number of species
$\frac{N_{\mathsf{out}}}{N}\approx\Phi\left(\frac{1/{\lambda^{*}}-N{\lambda^{*}}}{\sqrt{N}}\right)\,.$ (S4.35)
@@
$\sum_{n=1}^{1/{\lambda^{*}}}\frac{1}{n}e^{-(n-N{\lambda^{*}})^{2}/2N}\approx e^{-(1-N{\lambda^{*}})^{2}/2N}\sum_{n=1}^{1/{\lambda^{*}}}\frac{1}{n}\approx\sqrt{2\pi N}\mu(-\ln{\lambda^{*}})\,$, (S4.39)
where in the last approximation we evaluated the (slow-changing) exponential term in the range $1\leq n\leq 1/{\lambda^{*}}$ as $e^{-(1-N{\lambda^{*}})^{2}/2N}\approx e^{-N{\lambda^{*}}^{2}/2}\approx\sqrt{2\pi N}\mu$, and then plugged in our estimation of $\lambda^{*}$} in the low-migration regime (Eq. (12) in the main text). Combining the above gives (25) as reported in the main text.
@@ 5.1 Maximum abundances distribution
As described above, our model can be reduced to a birth-death process with transition probabilities $b_{n}$ and $d_{n}$, Eqs. (S1.6)-(S1.7). We are also using the approximation $\lambda(\mathbf{n})\approx{\lambda^{*}}$. Then, the probability $u_{n}(\ell)$ must obey two boundary conditions:
@@ 5.2 Bimodality of $Q_{m}$
$0=\frac{d}{dm}Q_{m}\approx-\frac{d^{2}}{dm^{2}}u_{1}(m)\propto-2\left(\frac{d}{dm}v_{m}\right)^{2}+v_{m}\frac{d^{2}}{dm^{2}}v_{m}\,$, (S5.51)
where, for notational convenience, we introduced $v_{m}=1/u_{1}(m)$ in the right side of Eq. (S5.49). Considering that equation, it can be shown that the derivatives of $v_{m}$ become very small at $m=N{\lambda^{*}}$, where the minimum of the exponent’s argument is reached. Linearizing the right side of Eq. (S5.51) around $m\approx N{\lambda^{*}}$ and then solving gives
@@
In Eq. (S5.57), we show that $\sqrt{2N}\mathscr{D}(x_{1})\approx-1/{\lambda^{*}}$. Since ${\lambda^{*}}\gg 1/\sqrt{N}$ in the low-migration regime, the denominator in the last term in Eq. (S5.52) is exponentially large, allowing us to ignore this whole term. We also have $N{\lambda^{*}}\gg 1/2$, thus dropping all sub-dominant terms gives the approximation
@@ 5.3 Infiltration probability
Here we derive the infiltration probability to the dynamical core, $\beta$. We will use the asymptotic expansion of the Dawson function [59, Eq. 42:6:4],
@@ 6 Residence times
In this section, we calculate the mean residence time of species whose trajectories reach the dynamical core, as defined in Section 5 above. We also argue that the cooperator core obtained in our study of the steady-state abundance distribution and the dynamical core are effectively equivalent.
At stationarity, let us denote by $T_{n}$ the mean first passage time (MFPT) from a species with initial abundance $n$ to reach extinction (at abundance 0). We observe that a core species will spend a long time at high abundance values; hence, we assume that the mean residence of a species whose trajectory belongs to the dynamical core can be approximated by the MFPT from $N{\lambda^{*}}$ to extinction,
@@
$V_{n}=U_{n-1}V_{1}-W_{n-1},\quad\text{with}\quad U_{m-1}:=\prod_{k=1}^{m-1}\frac{d_{k}}{b_{k}}\quad\text{and}\quad W_{m-1}=\sum_{k=1}^{m-1}\frac{1}{b_{k}}\prod_{j=k+1}^{m-1}\frac{d_{j}}{b_{j}}.$ (S6.61)
@@
$W_{m-1}=\sum_{k=1}^{m-1}\frac{1}{b_{k}}\prod_{j=k+1}^{m-1}\frac{d_{j}}{b_{j}}=U_{m-1}\sum_{k=1}^{m-1}\frac{1}{b_{k}}\prod_{j=1}^{k}\frac{b_{j}}{d_{j}}.$ (S6.63)
The expression for $W_{m-1}$ is related to the steady-state distribution $P_{k}$ because
@@
therefore we may rewrite $W_{m-1}$ as
$W_{m-1}=\frac{U_{m-1}}{P_{1}d_{1}}\sum_{k=1}^{m-1}P_{k}=\frac{U_{m-1}}{P_{1}d_{1}}C(m-1).$ (S6.65)
@@
The expected number of species in steady state, $R^{*}$, satisfies the balance equation between outflow due to extinction and inflow due to migration:
@@
Next, we observe that in the low-migration regime, $P_{k}$ is well-approximated as a Gaussian distribution with mean $N{\lambda^{*}}$ and standard deviation $\sqrt{N}$ (see Eq. (10) in the main text). Hence, $C$ is approximated as the cumulative density function of this Gaussian distribution, so $1-C(m-1)\approx 1$ for small $m$ and $1-C(m-1)\approx 1/2$ for $m\approx N{\lambda^{*}}$. On the other hand, for our range of $m\in[1,N{\lambda^{*}}]\cap\mathbb{N}$, $U_{m-1}$ has a local maximum at $m=1$ and, at small values of $m$, decays exponentially with rate $1/{\lambda^{*}}$, see Eq. (S6.62). Therefore, in the low-migration regime where $N{\lambda^{*}}^{2}\gg 1$, one can verify that as $m$ increases, $U_{m-1}$ decays essentially to 0 while 1C(m1)1-C(m-1) remains very close to 1. This allows us to approximate:
@@
The right-hand side corresponds to the expression for $1/u_{1}(N{\lambda^{*}})$ that we approximated in (S5.49). Using our approximation of $\beta\approx u_{1}(N{\lambda^{*}})$ (S5.55), we combine with (S6.72) to give
$T_{N{\lambda^{*}}}\approx\frac{R^{*}}{\mu\beta}.$ (S6.74)